CEBAF Operations Summaries
Recent Log Entries
- COMMENT ON Hall A Berm Improvements
- COMMENT ON 2026 SAM LCW Controls Calibrations
- iocsoftnl08rf restored
- iocsoftnl08rf restored
- COMMENT ON EA / WA LCW Systems - Install pH sensor
- COMMENT ON BE-CH-1&2 PM BEAMS chillers E&W Arc
- APPROVED Wien magnet work-CEBAF/Injector
- COMMENT ON 26/27 SAM Power Supply Flushing
- COMMENT ON Upgrade 1L08 to a C75 zone.
- Survey and Alignment work summary
- COMMENT ON Disconnect, Remove 20 Amp Trim Power Supplies For Bench Repair
- COMPLETED Disconnect Injector Wien Magnets MWF1I04 and MWF1I06
- iocr999rffcc1 restored
- COMMENT ON Core drlling for CEBAF Laser Room Cable Passthrough
- Follow-up Re: Ramped VWien+HWien off










